OpenStudy (soahelp):

Can somebody see if i'm correct on my one answer i will give a medal Which of the following is equivalent to (mn)3? 3mn m3n mn3 i think its this one m3n3

TheSmartOne (thesmartone):

uhhh is that \(\sf\large (mn)^3=?\)

TheSmartOne (thesmartone):

@SoaHelp

OpenStudy (alexandervonhumboldt2):

(ab)^3=a^3*b^3 use this property

OpenStudy (jango_in_dtown):

you are correct

OpenStudy (soahelp):

yes

TheSmartOne (thesmartone):

use the property \(\sf\large (mn)^x = m^xn^x\)

OpenStudy (soahelp):

the question its asking is if i use \[mn ^{3}\] will it equal the same as \[\left( mn \right)^{3}\]

OpenStudy (alexandervonhumboldt2):

no

OpenStudy (soahelp):

y not

OpenStudy (alexandervonhumboldt2):

(mn)^3=m^3*n^3

OpenStudy (alexandervonhumboldt2):

using the propetry (mn)^x=m^x*n^x

OpenStudy (soahelp):

one sec brb

OpenStudy (soahelp):

can you give me an example with it please

OpenStudy (soahelp):

@AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

(3*4)^9=3^9*4^9

OpenStudy (soahelp):

@AlexandervonHumboldt2 your going to kill me can you explain it to me

OpenStudy (soahelp):

like fill it out